OpenStudy (anonymous):

Given the quadratic function f(x) = 16x2 – 64x + 64, find a value of x such that f(x) = 16

OpenStudy (anonymous):

if f(x) = 16, then: 16x² - 64x + 64 = 16 simplify by dividing through by 16: x² - 4x + 4 = 1 subtract 1 from both sides: x² - 4x + 3 = 0 factor: (x - 3)(x - 1) = 0 solve: x = 3 or x = 1 There are two values of x such that f(x) = 16: f(3) = 16 and f(1) = 16

OpenStudy (anonymous):